Adult tickets for a play cost $16 and child tickets cost $6. If there were 25 people at a performance and the theater collected $260 from ticket sales, how many adults and how many children attended the play?
Answer by Cromlix(4381)

Children = x
Adults = y
x + y = 25...1
6x + 16y = 260...2
Multiply (1) by 6
6x + 6y = 150...3
6x + 16y = 260...2
Subtract (3) from (2)
10y = 110
y = 11.
Substitute y = 11 in (1)
x + y = 25
x + 11 = 25
x = 14.
14 children and 11 adults attended
the play.
Hope this helps.
:-)